OpenStudy (anonymous):

Lim sqrt (2c-a) -a/a-c x -> ∞

OpenStudy (anonymous):

\[\lim x ->\infty \left(\begin{matrix}\sqrt{(2c-a)} -a\\ a-c\end{matrix}\right)\]

OpenStudy (dan815):